# Convective Flows in Sheared Packings of Spherical Particles

### v1

@@ Abstract
We have demonstrated direct experimental evidence of geometry-driven convective flow in sheared granular packings of spherical particles. Using tracer layers and tomographic reconstruction, we reveal a transition from symmetric counter-rotating convection rolls at low filling heights to a dominant single-roll pattern at higher fillings. This transition correlates with increasing asymmetry in the radial shear profile, consistent with a shear-driven convection mechanism. These findings highlight the intrinsic ability of granular systems to sustain bulk transport through internal flow structures, even in the absence of particle anisotropy or boundary effects. Such geometry-induced convection can play a crucial role in mixing, segregation, and particle migration in granular media, with direct relevance to natural processes and industrial applications.

### v2

@@ Abstract
We have presented experimental evidence of geometry-driven convective flow in sheared granular packings of spherical particles. Using tracer layers and tomographic reconstruction, we reveal a transition from symmetric counter-rotating convection rolls at low filling heights to a dominant single-roll pattern at higher fillings. This transition correlates with increasing asymmetry in the radial shear profile, consistent with a shear-driven convection mechanism. Our work provides direct experimental evidence that secondary convection rolls can arise in the bulk of spherical granular materials, even in the absence of particle-shape anisotropy or boundary-driven effects, extending beyond previous simulation-based studies. Such geometry-induced convection can play a crucial role in mixing, segregation, and particle migration in granular media, with direct relevance to natural processes and industrial applications.